### v1

@@ Appendix E Testing $s$-sparsity
We will try to estimate the weight of each bucket within an additive error of $\tau/3$, where $\tau\geq\frac{1}{(p^{2}s)^{p-1}}$. If the function $f$ is indeed $s$-sparse, then clearly the number of buckets with estimated weight more than $32\tau/$ 3 is at most $s$. On the other hand, we will show that if the function $f$ is $\epsilon$-far from any $s$-sparse Boolean function, then with high probability at least $(s+1)$ buckets will have estimated weight more than $32\tau/$ 3.
Thus, the main problem boils down to estimating the weight of a bucket by querying $f$ at a small number of points. Unfortunately, for arbitrary buckets the number of queries to be made can be a lot! To bypass this, the buckets are choosen carefully. The buckets corresponds to the cosets of $H^{\perp}$ where $H$ is a random subspace of $\mathbb{Z}_{p}^{n}$ of dimenstion, $t=\Theta(s^{2})$. For such kind of buckets we show that estimation of the weight can be done using a small number of samples.
E.1 The Algorithm
Algorithm 1 takes a Boolean valued function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ as input. It accepts with probability $\geq\frac{2}{3}$ if $f$ is $s$-sparse, and rejects with probability $\geq\frac{2}{3}$ if $f$ is $\epsilon$-far from any $s$-sparse Boolean valued function.
In lines 4 and 5 we pick a random vector space of $\mathbb{Z}_{p}^{n}$ of dimension $t$, spanned by $v_{1},\ldots,v_{t}$.
For each $b=(b_{1},\ldots,b_{t})$ in $\mathbb{Z}_{p}^{t}$, we define the bucket $C(b)$ as follows:
$C(b):=\{\chi\in\mathbb{Z}_{p}^{n}:\langle\chi,v_{i}\rangle=b_{i}\ \forall i\in\{1,2,\ldots,t\}\}.$
We perform a random shift $u\in\mathbb{Z}_{p}^{t}$.12) and rename the buckets $C(b)$ as $C_{u}(b)$. Observe that the buckets $C_{u}(b)$ gives a partition of $\mathbb{Z}_{p}^{n}$ and each $C_{u}(b)$ corresponds to a coset in $\mathbb{Z}_{p}^{n}/H^{\perp}$. Thus once we have picked the random vector space $H$ in lines 4 and 5 we have a partition of $\mathbb{Z}_{p}^{n}$ (and thus the Dual space of $\mathbb{Z}_{p}^{n}$) into $p^{t}$ buckets where each bucket is of the form $(r+H^{\perp})$ for $r\in\mathbb{Z}_{p}^{n}$.
@@ E.1 The Algorithm
Lemma E.2 (Estimating the wt function).
For all $r\in\mathbb{Z}_{p}^{n}$, the value $\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$ calculated in Line 9 satisfy the following:
$\Pr\left(\left|\mathsf{wt}(r+H^{\perp})-\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})\right|\leq\frac{\tau}{3}\right)\geq 1-\frac{1}{10p^{t}}.$
The proof of Lemma E.2 is given in Section E.2. From Lemma E.2 we observe that by union bound (since there are $p^{t}$ buckets) at the end of line 9 with probability at least $\frac{9}{10}$ for all $r\in\mathbb{Z}_{p}^{n}$ the value of $\mathsf{wt}(r+H^{\perp})$ is estimated within additive $\frac{\tau}{3}$. That is, with probability $\frac{9}{10}$
$\forall r\in\mathbb{Z}_{p}^{n},\ \left|\mathsf{wt}(r+H^{\perp})-\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})\right|\leq\frac{\tau}{3}\;.$ (E.1)
@@
Lemma E.3 (Completeness theorem).
If $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ is an $s-$sparse Boolean function, then with probability $\frac{2}{3}$, the test accepts.
Proof E.4.
@@
It thus follows from Equation E.1 that with probability $\frac{9}{10}$, for each bucket with zero weight the estimation is $\leq\frac{\tau}{3}$ and for bucket with non-zero weights the estimation is $\geq\frac{2\tau}{3}$. Thus, in line 12 the algorithm will output YES with probability at least $\frac{9}{10}$.
@@
Lemma E.5.
If $f:\mathbb{Z}_{p}^{n}\to\{+1,-1\}$ is $\epsilon$-far from any $s$-sparse Boolean function, where $\epsilon\leq\frac{1}{8(p^{2}s)^{p-1}}$, then with probability $\frac{9}{10}$ there are at least $s+1$ buckets (that is at least $s+1$ vectors $r$ in $\mathbb{Z}_{p}^{n}$) such that weights of these buckets are more that $\tau$.
The proof of Lemma E.5 is given in Section E.3. Using Lemma E.5 we have the soundness of our algorithm.
Lemma E.6.
If $f$ is $\epsilon$-far from any $s$-sparse Boolean function, where $\epsilon\leq\frac{1}{8(p^{2}s)^{p-1}}$, the test rejects with probability $\frac{2}{3}$.
Proof E.7.
Let $f$ be $\epsilon$-far from any $s$-sparse Boolean function. By Lemma E.5 with probability $\frac{9}{10}$ there are at least $(s+1)$ buckets with weight at least $\tau$.
From Equation E.1, with probability $\frac{9}{10}$ the estimations of all the bucket are indeed accurate and thus for each bucket with weight $\geq\tau$ the estimation is $\geq\frac{2\tau}{3}$. And so in line 14 the algorithm will output NO.
Finally we observe that the algorithm makes $O(M)=O(\log(p^{t})\cdot\frac{1}{\tau^{2}})=\mathsf{poly}((p^{2}s)^{p},\frac{1}{\epsilon})$ samples.
E.2 Proof of Lemma E.2
@@
Lemma E.8.
$\mathbb{E}_{x\in\mathbb{Z}_{p}^{n},z\in H}[\chi_{r}(z)f(x)f(x-z)]=\mathsf{wt}(r+H^{\perp})$.
Proof E.9 (Proof of Lemma E.8).
For the proof of this lemma we need the projection operator defined in Definition B.1. Recall given a subspace $H^{\perp}\subseteq\mathbb{Z}_{p}^{n}$ and a coset $r+H^{\perp}$, we define the projection operator $P_{r+H^{\perp}}f$ on functions $f:\mathbb{Z}_{p}^{n}\to\mathbb{R}$ as follows:
@@
Recall that $wt(r+H^{\perp})=\sum_{\alpha\in r+H^{\perp}}|\hat{f}(\alpha)|^{2}$ which is equal to $\sum_{\alpha\in\mathbb{Z}_{p}^{n}}|\widehat{P_{r+H^{\perp}}f}(\alpha)|^{2}$, by the definition of $\widehat{P_{r+H^{\perp}}f}$. By Parseval, we know that
$\sum_{\alpha\in\mathbb{Z}_{p}^{n}}|\widehat{P_{r+H^{\perp}}f}(\alpha)|^{2}=\mathbb{E}_{w\in\mathbb{Z}_{p}^{n}}[P_{r+H}f(w)\overline{P_{r+H}f(w)}].$
From Lemma B.2 we have $P_{r+H^{\perp}}f(x)=\mathbb{E}_{z\in H}[f(x-z)\chi_{r}(z)]$ and hence $\overline{P_{r+H}f(x)}=\mathbb{E}_{z\in H^{\perp}}[f(x-z)\overline{\chi_{r}(z)}]$. Thus,
$\mathbb{E}_{w\in\mathbb{Z}_{p}^{n}}[P_{r+H}f(w)\overline{P_{r+H}f(w)}]\;=\;\mathbb{E}_{w\in\mathbb{Z}_{p}^{n},z_{1},z_{2}\in H^{\perp}}[\chi_{r}(z_{1})f(w-z_{1})\overline{\chi_{r}(z_{2})}f(w-z_{2})]\;.$
Since, the conjugate of $w_{p}$ ($p$-th root of unity) is $w_{p}^{p-1}$, by definition of $\chi_{r}(\cdot)$, we have
@@
$wt(r+H)=\mathbb{E}_{x\in\mathbb{Z}_{p}^{n},z\in H^{\perp}}[\chi_{r}(z)f(x)f(x-z)]\;$
as desired.
We will be using the Hoeffding’s Inequality (Lemma A.22) for our proof of correctness.
Now we are ready to prove Lemma E.2.
Proof E.10 (Proof of Lemma E.2).
Define $\delta:=\frac{1}{10p^{t}}$. Define $Y_{i}:=\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$. Write $\chi_{r}(z_{i})=a_{r}(z_{i})+\iota b_{r}(z_{i})$, where $a_{r}(z_{i})$ and $b_{r}(z_{i})$ are real numbers. Furthermore, since $|\chi_{r}(z_{i})|=1$ (Lemma A.2 (4)), we have that $a_{r}(z_{i}),b_{r}(z_{i})\in[-1,1]$. Define, $Y_{1,i}:=a_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$, and $Y_{2,i}:=b_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$. Since $f$ is Boolean valued, we have that $Y_{1,i},Y_{2,i}\in[-1,1]$. By definition, $Y_{i}=Y_{1,i}+\iota Y_{2,i}$. Consider $Y_{1}:=\frac{\sum_{i\in[M]}Y_{1,i}}{M}$ and $Y_{2}:=\frac{\sum_{i\in[M]}Y_{2,i}}{M}$. Since the pairs $(x_{i},z_{i})$ are all drawn independently, we have that the random variables $\{Y_{1,i}\}_{i=1}^{k}$ are independent, and the random variables $\{Y_{2,i}\}_{i=1}^{k}$ are independent. By Hoeffding’s Inequality (Lemma A.22) we have that for each $i\in[2]$,
$\Pr\left[|Y_{i}-\mathbb{E}[Y_{i}]|\geq\frac{\tau}{\sqrt{2}}\right]$
[…]
where the last inequality follows by an appropriate choice of the constant hidden in the definition $M=O(\log(p^{t})\cdot\frac{1}{\tau^{2}})$ in Algorithm 1. Let $Y=Y_{1}+\iota Y_{2}=\frac{\sum_{i\in[n]}Y_{i}}{M}$. Note that, $(Y-\mathbb{E}[Y])^{2}=(Y_{1}-\mathbb{E}[Y_{1}])^{2}+(Y_{2}-\mathbb{E}[Y_{2}])^{2}$. Trivially, $|Y-\mathbb{E}[Y]|\geq\tau\implies\exists i\in[2]$, such that $|Y_{i}-\mathbb{E}[Y_{i}]|\geq\frac{\tau}{\sqrt{2}}$. Therefore, we have
@@
From Lemma E.8 we have that $\mathbb{E}[Y]=\mathsf{wt}(r+H^{\perp})$. Hence we have our lemma.
E.3 Proof of Lemma E.5
Before we go on to prove Lemma E.5 it is important to understand some structure of the partition of the dual set we obtain in Lines 5 and 9. Recall that after we select the subspace $H$ at random the dual set is partitioned into buckets $C(b)$, where,
$C(b)=\{\chi\in\mathbb{Z}_{p}^{n}:\langle\chi,v_{i}\rangle=b_{i}\ \forall i\in\{1,2,\ldots,t\}\},$
where $v_{i}$ are the rows of the matrix $A$.
@@
Definition E.12 (Random shift).
Let us pick a non-zero element $u\in\mathbb{Z}_{p}^{t}$. For each $b$, let us define $C(b+u)$ as $C_{u}(b)$. This is known as a random shift. It is easy to see that $C_{u}(b)$ is also a coset and is isomorphic to $C(b)$.
Recall the definition of covariance of two random variables (Definition A.19). We need the following theorem.
Theorem E.13.
@@
    For all $\chi\in\mathbb{Z}_{p}^{n}$ and for all $b\in\mathbb{Z}_{p}^{t}$, $\Pr_{u,A}[\chi\in C_{u}(b)]=\frac{1}{p^{t}}$.
@@
    Let $S\subseteq\mathbb{Z}_{p}^{n}$ be such that $|S|\leq s+1$. Then, if $t\geq 2\log_{p}s+\log_{p}(\frac{1}{\delta})$, all $\chi\in S$ belong to different buckets except with probability at most $\delta$.
@@
    For all $b\in\mathbb{Z}_{p}^{t}$ and for all distinct $\chi,\chi^{\prime}\in\widehat{\mathbb{Z}_{p}^{n}}$, $\mathsf{cov}(I_{\chi\to b+u},I_{\chi^{\prime}\to b+u})=0$.
Proof E.14.
@@
    The event $\chi\in C_{u}(b)$ is equivalent to the event $\forall i\in[t],\langle\chi,v_{i}\rangle=b_{i}+u_{i}$. For each $i\in[t]$, $\Pr[\langle\chi,v_{i}\rangle=b_{i}+u_{i}]=\Pr[u_{i}=\langle\chi,v_{i}\rangle-b_{i}]=1/p$. Since, $u_{1},u_{2},\ldots,u_{t}$ and $b_{1},\ldots,b_{t}$ are all independent, the statement follows.
@@
    Let $\chi$ and $\chi^{\prime}$ be two distinct characters in $\widehat{\mathbb{Z}_{p}^{n}}$. We have that
    $\Pr\left[\chi,\chi^{\prime}\text{ belong to the same bucket}\right]=\Pr\left[\forall i\in[t],\langle\chi-\chi^{\prime},v_{i}\rangle=0\right]=\frac{1}{p^{t}},$
@@
    Pr[all $\Pr[\text{all }\chi\in S\text{ belong to different buckets}]$ $\geq 1-\binom{s+1}{2}\frac{1}{p^{t}}$
    $\geq 1-s^{2}\frac{1}{p^{t}}$
    $\geq 1-s^{2}\frac{1}{p^{2\log_{p}s+\log_{p}(\frac{1}{\delta})}}$
@@
    For all $b$, $\Pr_{u}[b+u\neq 0]=\frac{p^{t}-1}{p^{t}}$. Now, let us look at the following cases to prove the theorem.
    Case 1: $\chi^{\prime}$ is a scalar multiple of $\chi$, and both of them are nonzero.
    In this case,
    $I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}=1$ if and only if

### v2

@@ 6 Testing $s$-sparsity
We will try to estimate the weight of each bucket within an additive error of $\tau/3$, where $\tau\geq\frac{1}{2^{\varphi(\Delta)}s_{f}^{\varphi(\Delta)}}$. If the function $f$ is indeed $s$-sparse, then clearly the number of buckets with estimated weight more than $2\tau/3$ is at most $s$. On the other hand, we will show that if the function $f$ is $\epsilon$-far from any $s$-sparse Boolean function, then with high probability at least $(s+1)$ buckets will have estimated weight more than $2\tau/3$.
Thus, the main problem boils down to estimating the weight of a bucket by querying $f$ at a small number of points. Unfortunately, for arbitrary buckets the number of queries to be made can be a lot! To bypass this, the buckets are choosen carefully. The buckets corresponds to the cosets of $H^{\perp}$ where $H$ is a random subspace of $\mathcal{G}$ of dimenstion, $t=\Theta(s^{2})$. For such kind of buckets we show that estimation of the weight can be done using a small number of samples.
Algorithm 1 Test Sparsity
1:Input: $s,\epsilon$, and query access to $f:\mathcal{G}\to\{-1,+1\}$.
3:Parameter setting: Set $t:=\lceil 2\log_{\Delta}s+\log_{\Delta}20\rceil+1,\tau:=\min(\frac{\epsilon^{2}}{40\Delta^{t}},\frac{1}{(\Delta^{2}s)^{\varphi(\Delta)}})$ $M:=O(\log(\Delta^{t})\cdot\frac{1}{\tau^{2}})$
4:Choose $v_{1},\dots,v_{t}$ independent elements uniformly at random from $\mathcal{G}$.
5:Let $H=\operatorname{Span}\{v_{1},\dots,v_{t}\}$
6:Pick $(z_{1},x_{1}),\dots,(z_{M},x_{M})$ uniformly and independently from $H\times\mathcal{G}$
7:Query $f(x_{1}),\dots,f(x_{M})$ and $f(x_{1}-z_{1}),\dots,f(x_{M}-z_{M})$
8:for For every $r\in\mathcal{G}$ do
10:if number of $r$ for which the estimate of $wt(r+H^{\perp})$ is $\geq\frac{2\tau}{3}$ is $\leq s$ then
11:  Output YES
12:else
13:  Output NO
6.1 The Algorithm
Algorithm 1 takes a Boolean valued function $f:\mathcal{G}\to\{-1,+1\}$ as input. It accepts with probability $\geq\frac{2}{3}$ if $f$ is $s$-sparse, and rejects with probability $\geq\frac{2}{3}$ if $f$ is $\epsilon$-far from any $s$-sparse Boolean valued function.
In lines 4 and 5 we pick a random vector space of $\mathcal{G}$ of dimension $t$, spanned by $v_{1},\ldots,v_{t}$.
For each $b=(b_{1},\ldots,b_{t})$ in $\mathbb{Z}_{\Delta}^{t}$, where $\Delta=\operatorname{LCM}\{p_{1}^{m_{1}},\ldots,p_{T}^{m_{T}}\}$, we define the bucket $C(b)$ as follows:
$C(b):=\{\chi\in\mathcal{G}:\langle\chi,v_{i}\rangle=b_{i}\ \forall i\in\{1,2,\ldots,t\}\}.$
We perform a random shift $u\in\mathbb{Z}_{\Delta}^{t}$6.9) and rename the buckets $C(b)$ as $C_{u}(b)$. Observe that the buckets $C_{u}(b)$ gives a partition of $\mathcal{G}$ and each $C_{u}(b)$ corresponds to a coset in $\mathcal{G}/H^{\perp}$. Thus once we have picked the random vector space $H$ in lines 4 and 5 we have a partition of $\mathcal{G}$ (and thus the Dual space of $\mathcal{G}$) into $\Delta^{t}$ buckets where each bucket is of the form $(r+H^{\perp})$ for $r\in\mathcal{G}$.
@@ 6.1 The Algorithm
Lemma 6.2 (Estimating the wt function).
For all $r\in\mathcal{G}$, the value $\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$ calculated in Line 9 satisfy the following:
$\Pr\left(\left|\mathsf{wt}(r+H^{\perp})-\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})\right|\leq\frac{\tau}{3}\right)\geq 1-\frac{1}{10\Delta^{t}}.$
The proof of Lemma 6.2 is given in Section 6.2. From Lemma 6.2 we observe that by union bound (since there are $\Delta^{t}$ buckets) at the end of line 9 with probability at least $\frac{9}{10}$ for all $r\in\mathcal{G}$ the value of $\mathsf{wt}(r+H^{\perp})$ is estimated within additive $\frac{\tau}{3}$. That is, with probability $\frac{9}{10}$
$\forall r\in\mathcal{G},\ \left|\mathsf{wt}(r+H^{\perp})-\frac{1}{M}\sum_{i=1}^{M}\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})\right|\leq\frac{\tau}{3}\;.$ (6.3)
@@
Lemma 6.4 (Completeness theorem).
If $f:\mathcal{G}\to\{-1,+1\}$ is an $s$-sparse Boolean function, then with probability $\frac{2}{3}$, the test accepts.
Proof.
@@
It thus follows from Eq. 6.3 that with probability $\frac{9}{10}$, for each bucket with zero weight the estimation is $\leq\frac{\tau}{3}$ and for bucket with non-zero weights the estimation is $\geq\frac{2\tau}{3}$. Thus, in line 11 the algorithm will output YES with probability at least $\frac{9}{10}$. ∎
@@
Lemma 6.5.
If $f:\mathcal{G}\to\{+1,-1\}$ is $\epsilon$-far from any $s$-sparse Boolean function, where $\epsilon\leq\frac{1}{8(2s)^{\varphi(\Delta)}}$, then with probability $\frac{9}{10}$ there are at least $s+1$ buckets (that is at least $s+1$ vectors $r$ in $\mathcal{G}$) such that weights of these buckets are more that $\tau$.
The proof of Lemma 6.5 is given in Section 6.3. Using Lemma 6.5 we have the soundness of our algorithm.
Lemma $6.$
If $f$ is $\epsilon$-far from any $s$-sparse Boolean function, where $\epsilon\leq\frac{1}{8(2s)^{\varphi(\Delta)}}$, the test rejects with probability $\frac{2}{3}$.
Proof.
Let $f$ be $\epsilon$-far from any $s$-sparse Boolean function. By Lemma 6.5 with probability $\frac{9}{10}$ there are at least $(s+1)$ buckets with weight at least $\tau$.
From Eq. 6.3, with probability $\frac{9}{10}$ the estimations of all the bucket are indeed accurate and thus for each bucket with weight $\geq\tau$ the estimation is $\geq\frac{2\tau}{3}$. And so in line 13 the algorithm will output NO. ∎
Finally we observe that the algorithm makes $O(M)=O(\log(\Delta^{t})\cdot\frac{1}{\tau^{2}})=\mathsf{poly}((2s)^{\varphi(\Delta)},\frac{1}{\epsilon})$ samples.
6.2 Proof of Lemma 6.2
@@
Lemma 6.7.
$\mathbb{E}_{x\in\mathcal{G},z\in H}[\chi_{r}(z)f(x)f(x-z)]=\mathsf{wt}(r+H^{\perp})$.
Proof of Lemma 6.7.
For the proof of this lemma we need the projection operator defined in Definition 3.1. Recall given a subgroup $H^{\perp}\subseteq\mathcal{G}$ and a coset $r+H^{\perp}$, we define the projection operator $P_{r+H^{\perp}}f$ on functions $f:\mathcal{G}\to\mathbb{R}$ as follows:
@@
Recall that $wt(r+H^{\perp})=\sum_{\alpha\in r+H^{\perp}}|\hat{f}(\alpha)|^{2}$ which is equal to $\sum_{\alpha\in\mathcal{G}}|\widehat{P_{r+H^{\perp}}f}(\alpha)|^{2}$, by the definition of $\widehat{P_{r+H^{\perp}}f}$. By Parseval, we know that
$\sum_{\alpha\in\mathcal{G}}|\widehat{P_{r+H^{\perp}}f}(\alpha)|^{2}=\mathbb{E}_{w\in\mathcal{G}}[P_{r+H}f(w)\overline{P_{r+H}f(w)}].$
From Lemma 3.2 we have $P_{r+H^{\perp}}f(x)=\mathbb{E}_{z\in H}[f(x-z)\chi_{r}(z)]$ and hence $\overline{P_{r+H}f(x)}=\mathbb{E}_{z\in H^{\perp}}[f(x-z)\overline{\chi_{r}(z)}]$. Thus,
$\mathbb{E}_{w\in\mathcal{G}}[P_{r+H}f(w)\overline{P_{r+H}f(w)}]\;=\;\mathbb{E}_{w\in\mathcal{G},z_{1},z_{2}\in H^{\perp}}[\chi_{r}(z_{1})f(w-z_{1})\overline{\chi_{r}(z_{2})}f(w-z_{2})]\;.$
Since, the conjugate of $w_{\Delta}$ ($\Delta$-th root of unity) is $w_{\Delta}^{\Delta-1}$, by definition of $\chi_{r}(\cdot)$, we have
@@
$wt(r+H)=\mathbb{E}_{x\in\mathcal{G},z\in H^{\perp}}[\chi_{r}(z)f(x)f(x-z)]\;$
as desired. ∎
We will be using the Hoeffding’s Inequality (Lemma 2.31) for our proof of correctness.
Now we are ready to prove Lemma 6.2.
Proof of Lemma 6.2.
Define $\delta:=\frac{1}{10\Delta^{t}}$. Define $Y_{i}:=\chi_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$. Write $\chi_{r}(z_{i})=a_{r}(z_{i})+\iota b_{r}(z_{i})$, where $a_{r}(z_{i})$ and $b_{r}(z_{i})$ are real numbers. Furthermore, since $|\chi_{r}(z_{i})|=1$ (Lemma 2.4 (4)), we have that $a_{r}(z_{i}),b_{r}(z_{i})\in[-1,1]$. Define, $Y_{1,i}:=a_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$, and $Y_{2,i}:=b_{r}(z_{i})f(x_{i})f(x_{i}-z_{i})$. Since $f$ is Boolean valued, we have that $Y_{1,i},Y_{2,i}\in[-1,1]$. By definition, $Y_{i}=Y_{1,i}+\iota Y_{2,i}$. Consider $Y_{1}:=\frac{\sum_{i\in[M]}Y_{1,i}}{M}$ and $Y_{2}:=\frac{\sum_{i\in[M]}Y_{2,i}}{M}$. Since the pairs $(x_{i},z_{i})$ are all drawn independently, we have that the random variables $\{Y_{1,i}\}_{i=1}^{k}$ are independent, and the random variables $\{Y_{2,i}\}_{i=1}^{k}$ are independent. By Hoeffding’s Inequality (Lemma 2.31) we have that for each $i\in[2]$,
$\Pr\left[\Big{|}Y_{i}-\mathbb{E}[Y_{i}]\Big{|}\geq\frac{\tau}{\sqrt{2}}\right]\leq\;$ $\Pr\left[\Big{|}\sum_{j\in[M]}Y_{i,j}-\mathbb{E}\Big{[}\sum_{j\in[M]}Y_{i,j}\Big{]}\Big{|}\geq\frac{\tau M}{\sqrt{2}}\right]$
$\leq 2\exp\left(-\frac{\tau^{2}M}{4}\right)$
$\leq\frac{\delta}{2}\;,$
where the last inequality follows by an appropriate choice of the constant hidden in the definition $M=O(\log(\Delta^{t})\cdot\frac{1}{\tau^{2}})$ in Algorithm 1. Let $Y=Y_{1}+\iota Y_{2}=\frac{\sum_{i\in[n]}Y_{i}}{M}$. Note that, $(Y-\mathbb{E}[Y])^{2}=(Y_{1}-\mathbb{E}[Y_{1}])^{2}+(Y_{2}-\mathbb{E}[Y_{2}])^{2}$. Trivially, $|Y-\mathbb{E}[Y]|\geq\tau\implies\exists i\in[2]$, such that $|Y_{i}-\mathbb{E}[Y_{i}]|\geq\frac{\tau}{\sqrt{2}}$. Therefore, we have
@@
From Lemma 6.7 we have that $\mathbb{E}[Y]=\mathsf{wt}(r+H^{\perp})$. Hence we have our lemma. ∎
6.3 Proof of Lemma 6.5
Before we go on to prove Lemma 6.5 it is important to understand some structure of the partition of the dual set we obtain in Lines 5 and 9. Recall that after we select the subspace $H$ at random the dual set is partitioned into buckets $C(b)$, where,
$C(b)=\{\chi\in\mathcal{G}:\langle\chi,v_{i}\rangle=b_{i}\ \forall i\in\{1,2,\ldots,t\}\},$
where $v_{i}$ are the rows of the matrix AA.
@@
Definition 6.9 (Random shift).
Let us pick a non-zero element $u\in\mathbb{Z}_{\Delta}^{t}$. For each $b$, let us define $C(b+u)$ as $C_{u}(b)$. This is known as a random shift. It is easy to see that $C_{u}(b)$ is also a coset and is isomorphic to $C(b)$.
Recall the definition of covariance of two random variables (Definition 2.28). We need the following theorem.
Theorem 6.10.
@@
    For all $\chi\in\mathcal{G}$ and for all $b\in\mathbb{Z}_{\Delta}^{t}$, $\Pr_{u,A}[\chi\in C_{u}(b)]=\frac{1}{\Delta^{t}}$.
@@
    Let $S\subseteq\mathcal{G}$ be such that $|S|\leq s+1$. Then, if $t\geq 2\log_{\Delta}s+\log_{\Delta}(\frac{1}{\delta})$, all $\chi\in S$ belong to different buckets except with probability at most $\delta$.
@@
    For all $b\in\mathbb{Z}_{\Delta}^{t}$ and for all distinct $\chi,\chi^{\prime}\in\widehat{\mathcal{G}}$, $\mathsf{cov}(I_{\chi\to b+u},I_{\chi^{\prime}\to b+u})=0$.
Proof.
@@
    The event $\chi\in C_{u}(b)$ is equivalent to the event $\forall i\in[t],\langle\chi,v_{i}\rangle=b_{i}+u_{i}$. For each $i\in[t]$, $\Pr[\langle\chi,v_{i}\rangle=b_{i}+u_{i}]=\Pr[u_{i}=\langle\chi,v_{i}\rangle-b_{i}]=1/\Delta$. Since, $u_{1},u_{2},\ldots,u_{t}$ and $b_{1},\ldots,b_{t}$ are all independent, the statement follows.
@@
    Let $\chi$ and $\chi^{\prime}$ be two distinct characters in $\widehat{\mathcal{G}}$. Also note that the number of buckets is $\Delta^{t}$. We have that
    $\Pr\left[\chi,\chi^{\prime}\text{ belong to the same bucket}\right]=\Pr\left[\forall i\in[t],\langle\chi-\chi^{\prime},v_{i}\rangle=0\right]=\frac{1}{\Delta^{t}},$
@@
    $\Pr[\text{all }\chi\in S\text{ belong to different buckets}]$ $\geq 1-\binom{s+1}{2}\frac{1}{\Delta^{t}}$
    $\geq 1-s^{2}\frac{1}{\Delta^{t}}$
    $\geq 1-s^{2}\frac{1}{\Delta^{2\log_{\Delta}s+\log_{\Delta}(\frac{1}{\delta})}}$
@@
    For all $b$, $\Pr_{u}[b+u\neq 0]=\frac{\Delta^{t}-1}{\Delta^{t}}$. Now, let us look at the following cases to prove the theorem.
    Case 1: $\chi^{\prime}$ is a multiple of $\chi$, and both of them are nonzero.
    In this case, there exists a $\lambda\in\mathbb{Z}_{\Delta}$ such that $\chi^{\prime}=\lambda\chi$. Then,
    $\langle\chi^{\prime},v_{i}\rangle=\lambda\langle\chi,v_{i}\rangle\ \forall i\in[t]$
    $\Rightarrow(b_{i}+u_{i})(\lambda-1)=0\ \forall i\in[t]$
    $\Rightarrow(b+u)(\lambda-1)=0.$
    We have the following to subcases.
    Subcase 1: ($\lambda-1$) is invertible in $\mathbb{Z}_{\Delta}$.
    In this case, $I_{\chi\to b+u}\cdot I_{\chi^{\prime}\to b+u}=1$ if and only if